I need help with this problem, during the past year, ms.nguyen invested $9,000. Some of her money was invested in certificates of deposits paying 4% annually and the rest was invested in stocks paying 7.5% annually. She received a total of $500 from these investments for the year. How much did she invest in the certificates of deposit paying 4%?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 7.50% per annum ------------ Amount invested = y =
9000
Interest----- 500

Part I 4.00% per annum ---x
Part II 7.50% per annum ---y
Total investment
x + 1 y= 9000 -------------1
Interest on both investments
4.00% x + 7.50% y= 500
Multiply by 100
4 x + 7.5 y= 50000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -36000.00
Add this to (2)
0 x 3.5 y= 14000
divide by 3.5
y = 4000
Part I 4.00% $ 5000
Part II 7.50% $ 4000

CHECK
5000 --------- 4.00% ------- 200.00
4000 ------------- 7.50% ------- 300.00
Total -------------------- 500.00

m.ananth@hotmail.ca